OpenStudy (anonymous):

9i-(13+7i) Add or Subtract.

OpenStudy (anonymous):

9i-(13+7i) =9i+7i-13 =16i-13

OpenStudy (tkhunny):

Try again. That is not correct.

OpenStudy (anonymous):

13+7i =20i 9i-20i =-11i

OpenStudy (tkhunny):

Ummm... No. 9i - (13 + 7i) 9i - 13 - 7i 9i - 7i - 13 2i - 13 Combine LIKE terms only!